Thursday, December 9, 2010

Nightlog 2010-12-08

SA: David
SO: Patrick
Others: Simon, Martin initially, then Peter and Charl.


Successfully got both B-CAM (Big Apogee camera) and VI-CAM (small Apogee
camera) running together. Control of B-CAM now done on newly constructed
(thanks Hamish & Simon) BCAM PC, which can (in principle, once we know
how) run the FLI focuser, independent of tel. focus.

Current focus offset between the two cameras is ~1.5mm.

Had tracker failure (abnormal end) when hexapods flat-lined & radio
plugs pulled, plus the good old brakes release error, previously
attributed to interference and cured. Also seemed to be a problem with
overheating of anti-gravity motor, which was in excess of 30 deg.
Tracker recovered by Peter & Charl, who also bled the glycol to bring
the temp down.

Lots of fiddling to get a decent focus on VI-CAM, but not really a great
success for some reason TBD.

Some RSS grating flexure tests done during tracker fix.


No comments:

Post a Comment